OpenStudy (anonymous):

Use the quadratic formula and a calculator to find all real solutions, correct to three decimals. (If there is no real solution, enter NO REAL SOLUTION.) x2 − 0.017x − 0.064 = 0

OpenStudy (anonymous):

The quadratic formula is : \[x =( -b +\sqrt{b ^{2}-4ac} )/2a\] \[x =( -b -\sqrt{b ^{2}-4ac} )/2a\] where a is the coefficient of x^2, b the coeffecient of x and c the constant term. a = 1 b = 0.017 c = 0.064 Plug those values into the equations to find your answer. :)